OpenStudy (anonymous):

-{[2a+3(a-b)]-[4a-(b-3)]}

OpenStudy (angela210793):

-[(2a+3a-3b)-(4a-b+3)]=-[5a-3b-4a+b-3]=-a+2b+3

OpenStudy (anonymous):

{-a+2 b+3}